Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
magazzu - 6:58 Monday 10 June 2019 (46065) Print this report
Operator Report - Night Shift 9 June 2019

ITF found in locking acquisition, LOW_NOISE_3_SQZ reached at 21:15 UTC. From remote M. Mantovani enabled the BS AA at around 23:20 UTC, Science Mode started at 21:26 UTC.
From around 0:40 UTC to 3:20 UTC the IMC_ref camera showed a visible oscillation of the beam. Also the power related to the Input Mode Cleaner started to decrease, but the Horizon seemed to remain unaffected ( see attached plot #1 ). The oscillation disappeared after the unlock.
ITF unlocked at 3:20 UTC ( see attached plot #2 ), I was able to reach LOW_NOISE_3_SQZ at 4:13 UTC but unlocked a minute later.
Relocking in progress.

Guard Tour
22:33 UTC -23:06 UTC
0:32 UTC - 1:08 UTC
2:30 UTC - 3:02 UTC
4:11 UTC - 4:41 UTC

Pending
Contact Sentenac for UpsServer failure (No data for UPS server, TubePumps, and TubeStations).

Images attached to this report
Comments to this report:
mantovani - 8:47 Monday 10 June 2019 (46066) Print this report

The ITF was not reaching the Dark fringe state. It was due to a ~160 Hz oscillation in the SSFS. I fixed this by increasing the crossing between MC and SSFS (from 1.2 to 1.5, to be checked if it needs to come back). Then I've restored the BS angular control in full bandwidth disabled by Nicola to make the system more robust.

It is worth to notice that once the ITF could not make the transition from locking in Dark fringe to locked in Dark fringe (there was not any evident issue, may be the automation was stuck). 

It was solved by going in down and back in dark fringe

Search
×

Warning

×